OpenStudy (anonymous):

How do you find the directrix of a parabola? Like if you had y=x^2/32? Thanks!

OpenStudy (mimi_x3):

\[ y = \frac{x^2}{32} => 32y = x^2 => x^2 = 32y\]

OpenStudy (mimi_x3):

looks like you have to find the vertex first; then the directrix..

OpenStudy (lgbasallote):

^genius

OpenStudy (anonymous):

It told me to find the equation, which is what I thought the y=x^2/32 was. Because it is supposed to be y=? I don't think I phrased my first question right..

OpenStudy (mimi_x3):

a parabola is: \((x-h)^2 = 4A(y-k)\)

OpenStudy (mimi_x3):

What is exactly the question?

OpenStudy (anonymous):

Find the equation of the directrix of the parabola y=x^2/32

OpenStudy (anonymous):

Then it says the equation is y=__blank__

OpenStudy (mimi_x3):

Ok, are you able to find the vertex or the focal length?

OpenStudy (anonymous):

I honestly don't know what that means... Like I know what the vertex is, but focal length goes over my head.

OpenStudy (mimi_x3):

Tell what is the vertex for your question?

OpenStudy (mimi_x3):

tell me*

OpenStudy (anonymous):

Honestly, I don't know.. Um. I know what it is, but I don't really know how to find it. Super embarrassing & I'm super sorry.. I'm not very good at math.. but I know one of the equations you need to use is like (1/4)px^2 or something

OpenStudy (mimi_x3):

okay agree that parabola is \(x^2 = 4ay\) ?

OpenStudy (anonymous):

Sure. :)

OpenStudy (mimi_x3):

and when the parabola is like that the vertex is \(0,0\)agree?

OpenStudy (anonymous):

Yes

OpenStudy (mimi_x3):

\(V(0,0)\)

OpenStudy (mimi_x3):

then the focus is \((0,a)\)

OpenStudy (anonymous):

Okay, but this is equation isn't in that form so we don't know exactly what a is, right?

OpenStudy (mimi_x3):

your equation is \(x^2=32y\)

OpenStudy (anonymous):

Wait, a=8

OpenStudy (mimi_x3):

yes!

OpenStudy (anonymous):

So the focal point thing is (0,8) right?

OpenStudy (mimi_x3):

that is the focus, the focal length is 8. now what about the directrix? the directrix is \(y=-a\)

OpenStudy (anonymous):

so y=-8.

OpenStudy (mimi_x3):

yess!

OpenStudy (anonymous):

So what equation do we put that into? into the x^2=ay?

OpenStudy (mimi_x3):

that is just the question of parabola. which is similar to your equation. that is why you used it to find the focal length. you have already finished the question; you aready found the directrix

OpenStudy (anonymous):

But I thought I had to find the equation? Like the equation for the directrix?

OpenStudy (mimi_x3):

the equation of the directrix is \(y=-a\) and you just found it \(y=-8\) what else do you want to find?

OpenStudy (anonymous):

Oh. Silly me! Nothing, I just didn't realize it was that simple..

OpenStudy (mimi_x3):

well it is not that hard :)

OpenStudy (anonymous):

Haha, I guess not ! & Thanks for your help! :) Really appreciate it!